# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 9.3. Settings
The following settings are supported by Figure 9.4.
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $\mathcal{O}=(O,(G_{t}:t\in V(O)))$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$, with torsos being denoted by $H_{t}$. Write $O=:t(0),t(1),\ldots,t(m),t(0)$ with indices in $\mathbb{Z}_{m+1}$. Assume that the representatives of the links appear in this enumeration in the order $o(C),o(B),o(D),o(A)$. For each node $t$ of $O$ let $U^{\prime}_{t}:=V(G_{t})$ and $W^{\prime}_{t}:=\bigcup\,\{\,V(G_{s}):s\in V(O-t)\,\}$. Let $(U_{t},W_{t})$ denote the right-left-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
@@
Assume the (Tutte-bagel setting). (9.12). Assume that the torso $H_{t(0)}$ of $\mathcal{O}$ is 3-connected; so $H_{t(0)}$ stems from the $AC$-banana. Let $\{x^{\prime}_{1},x^{\prime}_{2}\}$ be the adhesion set of $t(-1)t(0)$ and let $\{y^{\prime}_{1},y^{\prime}_{2}\}$ be the adhesion set of $t(0)t(1)$. Let $(U^{\prime},W^{\prime}):=(U^{\prime}_{t(0)},W^{\prime}_{t(0)})$ and $(U,W):=(U_{t(0)},W_{t(0)})$. Note that $S(U^{\prime},W^{\prime})=\{x^{\prime}_{1},x^{\prime}_{2},y^{\prime}_{1},y^{\prime}_{2}\}$. Let $x_{i}\in S(U,W)$ be equal or incident to $x^{\prime}_{i}$, and let $x^{W}_{i}$ be the unique vertex in $\hat{x}_{i}\cap W$. Define $y_{i}$ and $y^{W}_{i}$ similarly.
@@
Assume the (Tutte-bagel torso setting). (9.13). The vertex-set $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(0)\,t(1)$ if the torso $H_{t(1)}$ is 3-connected, and it is the adhesion-set at $t(1)\,t(2)$ otherwise. In the latter case, the torso at $t(2)$ is not a 4-cycle.
@@
The former case is straightforward, so it remains to consider the case where $H_{t(1)}$ is not 3-connected. If the torso $H_{t(1)}$ is a 4-cycle, then the $y_{i}$ are the free edges of $H_{t(1)}$, and so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)t(2)$. Hence we may assume that the torso $H_{t(1)}$ is a triangle. The next torso $H_{t(2)}$ cannot be 4-cycle by Lemma 9.4, so it is 3-connected or a triangle.
Assume first that $H_{t(2)}$ is 3-connected. Since $H_{t(2)}$ shares a vertex with $H_{t(1)}$, say $y^{\prime}_{1}$, the vertex-set of $H_{t(2)}$ must be included in $C$. Hence the $\geqslant 2$ neighbours of $y^{\prime}_{1}$ in the bag $G_{t(2)}$ lie in $W^{\prime}\smallsetminus U^{\prime}$. Thus $y^{W}_{1}=y^{\prime}_{1}$. Since $H_{t(1)}$ is a triangle, $y_{2}$ is its free edge, so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)\,t(2)$.
Otherwise it is a triangle, so the adhesion-graph at $t(1)\,t(2)$ is a $K_{2}$-subgraph of $G$. Then the $y^{\prime}_{i}$ incident to the free edge of the triangle $H_{t(1)}$ corresponds to an edge $y_{i}$. The $y^{\prime}_{j}$ that is the tip of $H_{t(1)}$ has $y^{W}_{i}$ as a neighbour and is incident to the free edge of $H_{t(2)}$, so $y^{\prime}_{j}=y_{j}$. Hence $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)\,t(2)$. ∎
@@
(i). Suppose for a contradiction that two triangle-torsos of $\mathcal{O}$ have the same tip. Say the torsos are $H_{r}$ and $H_{t}$, and $v$ is their shared tip. Let $e$ be the free edge of $H_{r}$, and let $e^{\prime}$ be the free edge of $H_{t}$. Then $\{v,e,e^{\prime}\}$ is a mixed-3-separator, contradicting 4-connectivity by Lemma 6.1.
(ii). First, assume that there is a node $p$ on $P$ other than $r,t$ with $v\in G_{p}$ such that the torso $H_{p}$ is a 4-cycle or 3-connected. That the adhesion-sets at $G_{p}$ are disjoint then contradicts that $v$ lies in all three bags $G_{r},G_{p},G_{t}$. Otherwise, all torsos of all nodes on $P$ are triangles. Let $e$ be the edge of $P$ incident with $r$, and let $e^{\prime}$ be the edge of $P$ incident with $t$. By (i), if $|P|\geqslant 4$ then the adhesion sets of $e$ and $e^{\prime}$ are disjoint, which contradicts that $v$ lies in all bags of nodes in $P$ by $P\subseteq O_{v}$. Otherwise $|P|\leqslant 3$ and we get the statement. ∎
@@ 9.4. N-Paths lemmata
Since the Tutte-banana $\mathcal{T}$ for $K$ has no 3-connected torsos, it also has no 4-cycles as torsos by Lemma 9.4. Hence all torsos of $\mathcal{T}$ are triangles; in particular, both links adjacent to $K$ are represented by triangle-torsos of $\mathcal{T}$. Again by Lemma 9.4, it follows that all adhesion-graphs of $\mathcal{T}$ are $K_{2}$-subgraphs of $G$. ∎
@@
[…]
Suppose second that $\mathcal{T}$ has no 3-connected torsos. Then all torsos of $\mathcal{T}$ are triangles, both the $A$-link and the $D$-link are represented by triangle-torsos, and $a_{1}a_{2}\in E(G)$ by Lemma 9.18. In particular, exactly one of $a_{1},a_{2}$ lies in the $A$-link, say it is $a_{1}$. Now $a_{2}$ cannot lie in the $D$-link, since otherwise some edge dangles from the $D$-link through the $A$-link which by the (Crossing Lemma). (4.1) implies that the corner $AD$ is potter, a contradiction. Hence the path $a_{1}a_{2}$ is as desired. ∎
@@
By Lemma 9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(i-1)t(i)$. Moreover, the torso $H_{t(i)}$ is not a 4-cycle. We consider two cases, supported by Figure 9.5.
Suppose first that the torso at $t(i)$ is 3-connected. Then we find two $y^{W}_{1}$$y^{W}_{2}$ paths $Q_{1},Q_{2}$ in the torso at $t(i)$ that avoid the torso-edge $y^{W}_{1}y^{W}_{2}$. Note that $G_{t(i)}\subseteq G[C]$. If neither $Q_{j}$ uses the other torso-edge $z_{1}z_{2}$, then we are done. So we may assume that $Q_{1}$, say, uses $z_{1}z_{2}$. Let $e=u_{1}u_{2}$ be the $BD$-edge for the $B$-link. Let $k\geqslant i$ be maximum such that $t(k)$ stems from the $BC$-Tutte-banana. By Lemma 9.16, we find two disjoint $\{z_{1},z_{2}\}$$\{u_{1},u_{2}\}$ paths $R_{1},R_{2}$ in $G_{t(i+1)}\cup\cdots\cup G_{t(k)}$, taking the trivial paths if $k=i$. We apply the (One-Path Lemma). (9.19) (with $B$ in place of $A$) to the $BD$-edge $u_{1}u_{2}$ for the $B$-link to find a $u_{1}$$u_{2}$ path $R_{3}$ in $G[B\cap D]$ that uses at most one vertex in the $D$-link. Moreover, if $BD$ is not potter, then we find $R_{3}$ so that it avoids any prescribed vertex $d$ in the $D$-link. Then $P_{1}:=(Q_{1}-z_{1}z_{2})\cup R_{1}\cup R_{2}\cup R_{3}$ and $P_{2}:=Q_{2}$ are as desired.
Suppose otherwise that the torso at $t(i)$ is a triangle. Then $i=2$ and the torso at $t(1)$ is not 3-connected, by Lemma 9.14. Since the torso at $t(2)$ is a triangle, the torso at $t(1)$ cannot be a 4-cycle by Lemma 9.4, so it must be a triangle as well. Then the adhesion-graph of $t(1)t(2)$ is a $K_{2}$-subgraph of $G$ with vertices $y^{W}_{1},y^{W}_{2}$. Then we let $P_{2}:=Q_{2}:=y^{W}_{1}y^{W}_{2}$ and construct $P_{1}:=R_{1}\cup R_{2}\cup R_{3}$ as in the first case with the (One-Path Lemma). (9.19). ∎
Figure 9.5. Two paths found by Lemma 9.20. The vertices $y^{W}_{i}$ are blue. The path $P_{2}=Q_{2}$ is red. The path $P_{1}$ is green. The green vertices in the $B$-link are $u_{1},u_{2}$.
@@
Recall that not both corners $AD,BD$ are potter by 4.7.
@@ 9.5. Totally-nested tetra-separations from good bags
(ii) $\Rightarrow$ (i). Assume that $H_{t(0)}$ is good. If $|H_{t(0)}|\geqslant 6$, then $U^{\prime}\smallsetminus W^{\prime}$ has size $\geqslant 2$, so $(U,W)$ is a tetra-separation by 5.10. If $|H_{t(0)}|=5$ or $|H_{t(0)}|=4$, then $H_{t(0)}$ being good ensures via the neighbouring torsos that at least one or at least two vertices in $S(U^{\prime},W^{\prime})$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$, respectively, so $(U,W)$ is a tetra-separation by 5.10.
(i) $\Rightarrow$ (ii). Assume that $(U,W)$ fails to be a tetra-separation. Then by 5.10, the number of vertices in $S(U^{\prime},W^{\prime})$ with $\leqslant 1$ neighbour in $W^{\prime}\smallsetminus U^{\prime}$ fails to be $\geqslant 2-|U^{\prime}\smallsetminus W^{\prime}|$. Hence all sufficient conditions for $H_{t(0)}$ being good fail. ∎
@@
If $S(U,W)$ contains an edge, then $G-S(U,W)$ has exactly two components by 4-connectivity, so $(U,W)$ is half-connected. Otherwise $(U,W)=(U^{\prime},W^{\prime})$. We will show that $G[W\smallsetminus U]$ is connected. Let $\pi$ denote the adhesion-set of $\mathcal{O}$ that is spanned by the $BD$-edge for the $D$-link by Lemma 9.11.
We claim that every vertex $v\in W\smallsetminus U$ not in $\pi$ sends a fan of two paths to $\pi$ in $G[W\smallsetminus U]$. Indeed, $\pi$ together with $\{x^{\prime}_{1},x^{\prime}_{2}\}$, say, separates $v$ from $G_{t(0)}$. Since $G$ is 4-connected, there is a fan of four paths from $v$ to $\{x^{\prime}_{1},x^{\prime}_{2}\}\cup\pi$ by Lemma 6.2. The two paths ending in $\pi$ are as desired.
@@
If $H_{t(1)}$ is 3-connected, we have (i). Otherwise $H_{t(1)}$ is a 4-cycle or a triangle. Since $y_{2}$ is a vertex by assumption, $H_{t(1)}$ is a triangle and $y_{2}=y^{W}_{2}$ is its tip, by Lemma 9.14. Hence $y^{W}_{1}$ is incident to the free edge of $H_{t(1)}$. If $H_{t(2)}$ is 3-connected, then we have (i). Otherwise $H_{t(2)}$ is a triangle, by Lemma 9.4. By Lemma 9.15, $y^{W}_{1}$ is the tip of $H_{t(2)}$. Hence $y^{W}_{2}$ is not contained in the adhesion-set between $H_{t(2)}$ and $H_{t(3)}$, and in particular $y^{W}_{2}$ is not contained in $H_{t(3)}$. If $H_{t(3)}$ is 3-connected, we have (i). Otherwise $H_{t(3)}$ also is a triangle. By Lemma 9.15, $y^{W}_{1}$ is incident to the free edge of $H_{t(3)}$. By Lemma 9.4, the adhesion-graph between the triangle-torsos $H_{t(2)}$ and $H_{t(3)}$ is a $K_{2}$-subgraph of $G$. Hence the two edges of $H_{t(3)}$ incident to $y^{W}_{1}$ exist in $G$. Thus we have (ii). ∎
@@
Assume that (i) fails. Then there exists a cutvertex $v$ of $G_{t}$ that separates $a_{1}$ from $b_{1}$. Since $H_{t}$ is 3-connected and $G_{t}$ misses at most two of the torso-edges, $v$ is distinct from $a_{2}$ and $b_{2}$. Adding the edges $a_{1}a_{2}$ and $b_{1}b_{2}$ makes $v$ into the separator of a mixed-3-separation $(X,Y)$ of $H_{t}$ with $a_{1}\in X\smallsetminus Y$ and $b_{1}\in Y\smallsetminus X$. Since $S(X,Y)$ contains edges, $H_{t}-S(X,Y)$ has exactly two components. These components include the vertex-sets $\{a_{1},v,b_{2}\}$ and $\{a_{2},v,b_{1}\}$, respectively. Since $G$ is 4-connected, these inclusions are equalities, and $v$ has degree four in $H_{t}$. As $H_{t}$ is 3-connected, $H_{t}$ is a 4-wheel with rim $a_{1}a_{2}b_{1}b_{2}$. Since (i) fails by assumption, both torso-edges $a_{1}a_{2}$ and $b_{1}b_{2}$ must be missing from $G$. ∎
@@
Case $i<j$. This case is supported by Figure 9.6. If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$. Otherwise, by Lemma 9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges. In both cases, $F_{y}$ avoids $y_{2}$. Similarly, we obtain an $x^{W}_{1}$$\pi_{j-1}$ fan $F_{x}\subseteq G_{t(j)}$ avoiding $x_{2}$. By Lemma 9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$. Call these paths $P$ and $Q$. Then $F_{x}\cup P\cup Q\cup F_{y}$ is a union of two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
Figure 9.6. Case $i<j$ in the proof of Lemma 9.29. Here, $H_{t(i)}$ is 3-connected and $H_{t(j)}$ is a triangle. The unique vertex in $\pi_{i-1}-y^{W}_{1}$ is black.
Case $i\geqslant j$. This case is supported by Figure 9.7. We claim that the corner $BD$ is not potter, and assume for a contradiction that it is. Let $\pi$ denote the unique adhesion-set of the $BD$-Tutte-banana. Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by 4.7. Hence if $y_{1}$ is an edge, it neither dangles from the $C$-link through the $B$-link nor from the $B$-link through the $C$-link, by the (Crossing Lemma). (4.1). So $y^{W}_{1}\in C\smallsetminus D$, and hence $y^{W}_{1}\notin\pi$. Similarly, $x^{W}_{1}\notin\pi$. But this contradicts $i\geqslant j$, so the corner $BD$ is not potter.

### v2

@@ 9.3. Settings
The following settings are supported by Figure˜9.4.
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $\mathcal{O}=(O,(G_{t}:t\in V(O)))$ be the Tutte-bagel induced by $(A,B)$ and $(C,D)$, with torsos being denoted by $H_{t}$. Write $O=:t(0),t(1),\ldots,t(m),t(0)$ with indices in $\mathbb{Z}_{m+1}$. Assume that the representatives of the links appear in this enumeration in the order $o(C),o(B),o(D),o(A)$. For each node $t$ of $O$ let $U^{\prime}_{t}:=V(G_{t})$ and $W^{\prime}_{t}:=\bigcup\,\{\,V(G_{s}):s\in V(O-t)\,\}$. Let $(U_{t},W_{t})$ denote the right-left-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
@@
Assume the (Tutte-bagel setting). (9.12). Assume that the torso $H_{t(0)}$ of $\mathcal{O}$ is 3-connected; so $H_{t(0)}$ stems from the $AC$-banana. Let $\{x^{\prime}_{1},x^{\prime}_{2}\}$ be the adhesion set of t(1)t(0)t(-1)t(0) and let $\{y^{\prime}_{1},y^{\prime}_{2}\}$ be the adhesion set of $t(0)t(1)$. Let $(U^{\prime},W^{\prime}):=(U^{\prime}_{t(0)},W^{\prime}_{t(0)})$ and $(U,W):=(U_{t(0)},W_{t(0)})$. Note that $S(U^{\prime},W^{\prime})=\{x^{\prime}_{1},x^{\prime}_{2},y^{\prime}_{1},y^{\prime}_{2}\}$. Let $x_{i}\in S(U,W)$ be equal or incident to $x^{\prime}_{i}$, and let $x^{W}_{i}$ be the unique vertex in $\hat{x}_{i}\cap W$. Define $y_{i}$ and $y^{W}_{i}$ similarly.
@@
Assume the (Tutte-bagel torso setting). (9.13). The vertex-set $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(0)t(1)t(0)\,t(1) if the torso $H_{t(1)}$ is 3-connected, and it is the adhesion-set at t(1)t(2)t(1)\,t(2) otherwise. In the latter case, the torso at $t(2)$ is not a 4-cycle.
@@
The former case is straightforward, so it remains to consider the case where $H_{t(1)}$ is not 3-connected. If the torso $H_{t(1)}$ is a 4-cycle, then the $y_{i}$ are the free edges of $H_{t(1)}$, and so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at $t(1)t(2)$. Hence we may assume that the torso $H_{t(1)}$ is a triangle. The next torso $H_{t(2)}$ cannot be 4-cycle by Lemma˜9.4, so it is 3-connected or a triangle.
Assume first that $H_{t(2)}$ is 3-connected. Since $H_{t(2)}$ shares a vertex with $H_{t(1)}$, say $y^{\prime}_{1}$, the vertex-set of $H_{t(2)}$ must be included in $C$. Hence the $\geqslant 2$ neighbours of $y^{\prime}_{1}$ in the bag $G_{t(2)}$ lie in $W^{\prime}\smallsetminus U^{\prime}$. Thus $y^{W}_{1}=y^{\prime}_{1}$. Since $H_{t(1)}$ is a triangle, $y_{2}$ is its free edge, so $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(1)t(2)t(1)\,t(2).
Otherwise it is a triangle, so the adhesion-graph at t(1)t(2)t(1)\,t(2) is a $K_{2}$-subgraph of $G$. Then the $y^{\prime}_{i}$ incident to the free edge of the triangle $H_{t(1)}$ corresponds to an edge $y_{i}$. The $y^{\prime}_{j}$ that is the tip of $H_{t(1)}$ has $y^{W}_{i}$ as a neighbour and is incident to the free edge of $H_{t(2)}$, so $y^{\prime}_{j}=y_{j}$. Hence $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(1)t(2)t(1)\,t(2). ∎
@@
˜(i). Suppose for a contradiction that two triangle-torsos of $\mathcal{O}$ have the same tip. Say the torsos are $H_{r}$ and $H_{t}$, and $v$ is their shared tip. Let $e$ be the free edge of $H_{r}$, and let $e^{\prime}$ be the free edge of $H_{t}$. Then $\{v,e,e^{\prime}\}$ is a mixed-3-separator, contradicting 4-connectivity by Lemma˜6.1.
˜(ii). First, assume that there is a node $p$ on $P$ other than $r,t$ with $v\in G_{p}$ such that the torso $H_{p}$ is a 4-cycle or 3-connected. That the adhesion-sets at $G_{p}$ are disjoint then contradicts that $v$ lies in all three bags $G_{r},G_{p},G_{t}$. Otherwise, all torsos of all nodes on $P$ are triangles. Let $e$ be the edge of $P$ incident with $r$, and let $e^{\prime}$ be the edge of $P$ incident with $t$. By (i), if $|P|\geqslant 4$ then the adhesion sets of $e$ and $e^{\prime}$ are disjoint, which contradicts that $v$ lies in all bags of nodes in $P$ by $P\subseteq O_{v}$. Otherwise $|P|\leqslant 3$ and we get the statement. ∎
@@ 9.4. N-Paths lemmata
Since the Tutte-banana $\mathcal{T}$ for $K$ has no 3-connected torsos, it also has no 4-cycles as torsos by Lemma˜9.4. Hence all torsos of $\mathcal{T}$ are triangles; in particular, both links adjacent to $K$ are represented by triangle-torsos of $\mathcal{T}$. Again by Lemma˜9.4, it follows that all adhesion-graphs of $\mathcal{T}$ are $K_{2}$-subgraphs of $G$. ∎
@@
Suppose first that $\mathcal{T}$ has a 3-connected torso $H_{t}$. We find two disjoint $a_{i}$$H_{t}$ paths $P_{i}$ for $i=1,2$ in the $AD$-banana using Lemma˜9.16. We even get that both $P_{i}$ are subgraphs of $G$. Note that the ends of the $P_{i}$ in $H_{t}$ together form an adhesion-set. If $H_{t}$ is a $K_{4}$, then we link the ends of the $P_{i}$ in the bag $G_{t}$ by a path that avoids any prescribed vertex $d$ in the $D$-link and that uses no torso-edges. Otherwise $H_{t}$ is not a $K_{4}$, so we may use Lemma˜9.17 to link the ends of the $P_{i}$ in the bag $G_{t}$ with a path that avoids all vertices in the $D$-link and that uses no torso-edges.
Suppose second that $\mathcal{T}$ has no 3-connected torsos. Then all torsos of $\mathcal{T}$ are triangles, both the $A$-link and the $D$-link are represented by triangle-torsos, and $a_{1}a_{2}\in E(G)$ by Lemma˜9.18. In particular, exactly one of $a_{1},a_{2}$ lies in the $A$-link, say it is $a_{1}$. Now $a_{2}$ cannot lie in the $D$-link, since otherwise some edge dangles from the $D$-link through the $A$-link which by the (Crossing Lemma). (4.1) implies that the corner $AD$ is potter, a contradiction. Hence the path $a_{1}a_{2}$ is as desired. ∎
@@
By Lemma˜9.14, there is $i\in\{1,2\}$ such that $\{y^{W}_{1},y^{W}_{2}\}$ is the adhesion-set at t(i1)t(i)t(i-1)t(i). Moreover, the torso $H_{t(i)}$ is not a 4-cycle. We consider two cases, supported by Figure˜9.5.
Suppose first that the torso at $t(i)$ is 3-connected. Then we find two $y^{W}_{1}$$y^{W}_{2}$ paths $Q_{1},Q_{2}$ in the torso at $t(i)$ that avoid the torso-edge $y^{W}_{1}y^{W}_{2}$. Note that $G_{t(i)}\subseteq G[C]$. If neither $Q_{j}$ uses the other torso-edge $z_{1}z_{2}$, then we are done. So we may assume that $Q_{1}$, say, uses $z_{1}z_{2}$. Let $e=u_{1}u_{2}$ be the $BD$-edge for the $B$-link. Let $k\geqslant i$ be maximum such that $t(k)$ stems from the $BC$-Tutte-banana. By Lemma˜9.16, we find two disjoint $\{z_{1},z_{2}\}$$\{u_{1},u_{2}\}$ paths $R_{1},R_{2}$ in $G_{t(i+1)}\cup\cdots\cup G_{t(k)}$, taking the trivial paths if $k=i$. We apply the (One-Path Lemma). (9.19) (with $B$ in place of $A$) to the $BD$-edge $u_{1}u_{2}$ for the $B$-link to find a $u_{1}$$u_{2}$ path $R_{3}$ in $G[B\cap D]$ that uses at most one vertex in the $D$-link. Moreover, if $BD$ is not potter, then we find $R_{3}$ so that it avoids any prescribed vertex $d$ in the $D$-link. Then $P_{1}:=(Q_{1}-z_{1}z_{2})\cup R_{1}\cup R_{2}\cup R_{3}$ and $P_{2}:=Q_{2}$ are as desired.
Suppose otherwise that the torso at $t(i)$ is a triangle. Then $i=2$ and the torso at $t(1)$ is not 3-connected, by Lemma˜9.14. Since the torso at $t(2)$ is a triangle, the torso at $t(1)$ cannot be a 4-cycle by Lemma˜9.4, so it must be a triangle as well. Then the adhesion-graph of $t(1)t(2)$ is a $K_{2}$-subgraph of $G$ with vertices $y^{W}_{1},y^{W}_{2}$. Then we let $P_{2}:=Q_{2}:=y^{W}_{1}y^{W}_{2}$ and construct $P_{1}:=R_{1}\cup R_{2}\cup R_{3}$ as in the first case with the (One-Path Lemma). (9.19). ∎
Figure 9.5. Two paths found by Lemma˜9.20. The vertices $y^{W}_{i}$ are blue. The path $P_{2}=Q_{2}$ is red. The path $P_{1}$ is green. The green vertices in the $B$-link are $u_{1},u_{2}$.
@@
Recall that not both corners $AD,BD$ are potter by ˜4.7.
@@ 9.5. Totally-nested tetra-separations from good bags
˜(ii) $\Rightarrow$ ˜(i). Assume that $H_{t(0)}$ is good. If $|H_{t(0)}|\geqslant 6$, then $U^{\prime}\smallsetminus W^{\prime}$ has size $\geqslant 2$, so $(U,W)$ is a tetra-separation by ˜5.10. If $|H_{t(0)}|=5$ or $|H_{t(0)}|=4$, then $H_{t(0)}$ being good ensures via the neighbouring torsos that at least one or at least two vertices in $S(U^{\prime},W^{\prime})$ have exactly one neighbour in $W^{\prime}\smallsetminus U^{\prime}$, respectively, so $(U,W)$ is a tetra-separation by ˜5.10.
˜(i) $\Rightarrow$ ˜(ii). Assume that $(U,W)$ fails to be a tetra-separation. Then by ˜5.10, the number of vertices in $S(U^{\prime},W^{\prime})$ with $\leqslant 1$ neighbour in $W^{\prime}\smallsetminus U^{\prime}$ fails to be $\geqslant 2-|U^{\prime}\smallsetminus W^{\prime}|$. Hence all sufficient conditions for $H_{t(0)}$ being good fail. ∎
@@
If $S(U,W)$ contains an edge, then GS(U,W)G-S(U,W) has exactly two components by 4-connectivity, so $(U,W)$ is half-connected. Otherwise $(U,W)=(U^{\prime},W^{\prime})$. We will show that $G[W\smallsetminus U]$ is connected. Let $\pi$ denote the adhesion-set of $\mathcal{O}$ that is spanned by the $BD$-edge for the $D$-link by Lemma˜9.11.
We claim that every vertex $v\in W\smallsetminus U$ not in $\pi$ sends a fan of two paths to $\pi$ in $G[W\smallsetminus U]$. Indeed, $\pi$ together with $\{x^{\prime}_{1},x^{\prime}_{2}\}$, say, separates $v$ from $G_{t(0)}$. Since $G$ is 4-connected, there is a fan of four paths from $v$ to $\{x^{\prime}_{1},x^{\prime}_{2}\}\cup\pi$ by Lemma˜6.2. The two paths ending in $\pi$ are as desired.
@@
If $H_{t(1)}$ is 3-connected, we have ˜(i). Otherwise $H_{t(1)}$ is a 4-cycle or a triangle. Since $y_{2}$ is a vertex by assumption, $H_{t(1)}$ is a triangle and $y_{2}=y^{W}_{2}$ is its tip, by Lemma˜9.14. Hence $y^{W}_{1}$ is incident to the free edge of $H_{t(1)}$. If $H_{t(2)}$ is 3-connected, then we have ˜(i). Otherwise $H_{t(2)}$ is a triangle, by Lemma˜9.4. By Lemma˜9.15, $y^{W}_{1}$ is the tip of $H_{t(2)}$. Hence $y^{W}_{2}$ is not contained in the adhesion-set between $H_{t(2)}$ and $H_{t(3)}$, and in particular $y^{W}_{2}$ is not contained in $H_{t(3)}$. If $H_{t(3)}$ is 3-connected, we have ˜(i). Otherwise $H_{t(3)}$ also is a triangle. By Lemma˜9.15, $y^{W}_{1}$ is incident to the free edge of $H_{t(3)}$. By Lemma˜9.4, the adhesion-graph between the triangle-torsos $H_{t(2)}$ and $H_{t(3)}$ is a $K_{2}$-subgraph of $G$. Hence the two edges of $H_{t(3)}$ incident to $y^{W}_{1}$ exist in $G$. Thus we have ˜(ii). ∎
@@
Assume that ˜(i) fails. Then there exists a cutvertex $v$ of $G_{t}$ that separates $a_{1}$ from $b_{1}$. Since $H_{t}$ is 3-connected and $G_{t}$ misses at most two of the torso-edges, $v$ is distinct from $a_{2}$ and $b_{2}$. Adding the edges $a_{1}a_{2}$ and $b_{1}b_{2}$ makes $v$ into the separator of a mixed-3-separation $(X,Y)$ of $H_{t}$ with $a_{1}\in X\smallsetminus Y$ and $b_{1}\in Y\smallsetminus X$. Since $S(X,Y)$ contains edges, $H_{t}-S(X,Y)$ has exactly two components. These components include the vertex-sets $\{a_{1},v,b_{2}\}$ and $\{a_{2},v,b_{1}\}$, respectively. Since $G$ is 4-connected, these inclusions are equalities, and $v$ has degree four in $H_{t}$. As $H_{t}$ is 3-connected, $H_{t}$ is a 4-wheel with rim $a_{1}a_{2}b_{1}b_{2}$. Since ˜(i) fails by assumption, both torso-edges $a_{1}a_{2}$ and $b_{1}b_{2}$ must be missing from $G$. ∎
@@
Case $i<j$. This case is supported by Figure˜9.6. If $H_{t(i)}$ is 3-connected, we let $F_{y}$ by a $y^{W}_{1}$$\pi_{i}$ fan of two paths in $G_{t(i)}$ avoiding the vertex in $\pi_{i-1}-y^{W}_{1}$. Otherwise, by Lemma˜9.27, $H_{t(i)}$ is a triangle avoiding $y^{W}_{2}$ such that both edges incident to $y^{W}_{1}$ in $H_{t(i)}$ exist in $G$, and we let $F_{y}$ denote the $y^{W}_{1}$$\pi_{i}$ fan of two paths defined by these two edges. In both cases, $F_{y}$ avoids $y_{2}$. Similarly, we obtain an $x^{W}_{1}$$\pi_{j-1}$ fan $F_{x}\subseteq G_{t(j)}$ avoiding $x_{2}$. By Lemma˜9.16, we find two disjoint $\pi_{i}$$\pi_{j-1}$ paths in $G_{t(i+1)}\cup\cdots\cup G_{t(j-1)}$. Call these paths $P$ and $Q$. Then $F_{x}\cup P\cup Q\cup F_{y}$ is a union of two independent $x^{W}_{1}$$y^{W}_{1}$ paths in $G[W]-x_{2}-y_{2}$.
Figure 9.6. Case $i<j$ in the proof of Lemma˜9.29. Here, $H_{t(i)}$ is 3-connected and $H_{t(j)}$ is a triangle. The unique vertex in $\pi_{i-1}-y^{W}_{1}$ is black.
Case $i\geqslant j$. This case is supported by Figure˜9.7. We claim that the corner $BD$ is not potter, and assume for a contradiction that it is. Let $\pi$ denote the unique adhesion-set of the $BD$-Tutte-banana. Since $BD$ is potter, neither corner $BC$ nor $AD$ is potter by ˜4.7. Hence if $y_{1}$ is an edge, it neither dangles from the $C$-link through the $B$-link nor from the $B$-link through the $C$-link, by the (Crossing Lemma). (4.1). So $y^{W}_{1}\in C\smallsetminus D$, and hence $y^{W}_{1}\notin\pi$. Similarly, $x^{W}_{1}\notin\pi$. But this contradicts $i\geqslant j$, so the corner $BD$ is not potter.